PS/DR/Note 2002-184 (Min.)
Summary of the Meeting of the PS Management Board
held on 28.10.2002
Present : S. Baird, J. Boillot, M. Chanel, G. Daems, J.P.
Delahaye (Chairman), B. Frammery, G. Geschonke, J. Gruber, C. Hill, T. Kehrer
(Secretary), M. Lindroos, K.D. Metzmacher, J.P. Potier, J.P. Riunaud,(Chairman)
Invited: T. Eriksson (point 1)
Excused : R. Garoby, H. Haseroth, S.L. Neboux, E. Métral (Staff Association),
* * * * *
1. Report from the ADUC meeting
T. Eriksson gave a summary of the ADUC (AD Users' Committee) meeting which
was held on Friday 25.10.2002.
- Status report AD and
outlook 2003 (T.Eriksson)
- Status reports: AD experiments and outlook 2003
Some initial problems with the AD piquet operation
mode it can be difficult to find somebody at the other end of the
ATHENA: Antihydrogen detected in July/August, maybe laser
spectroscopy in 2003, no major changes to the experiment in 2003.
ATRAP: Detection of cold antihydrogen will be announced
soon. They are looking into spectroscopy for 2003. Still waiting for the
main magnet to be used in their second trap/detector.
ASACUSA: Some leftover measurements from 2002 done. Last
2 months spent on trap extraction development progress is slow.
Looking into antiprotonic helium spectroscopy for 2003. 60% of the work
that was planned until 2005 is done.
Comments on SPSC proposal P-234 Study of pbar-medical
applications (R.Landua, J.P. Riunaud)
Collaboration Pbar Medical Inc./UCLA medical/Aarhus
with 17 participants.
Relative biological effectiveness, peripheral damage
by pbar annihilation. Investigate cell survival as function of dose.
Requires : Use of AD DEM zone, 300 MeV/c extraction,
9 complete shifts for data taking and 1 week AD setting-up time.
Experiment will pay for required additional HW and manpower.
Preparation in 2002/early2003, data taking July-October
Presentation to SPSC 5/11/2002, then recommendation
to research board.
Comments from the existing users: Happy to have a 4:th
experiment but not too willing to give up any beamtime.
Discussion: "planning the AD future after 2005"
The users agree that something has to be proposed to
enlarge the user community and to the benefit of all users.
More pbars would be available for all users if the RFQD
was moved upstreams to supply all experiments via a switchyard and building
a low energy beam transport similar the what is planned in ISOLDE.
R.Hayano indicated that the ASACUSA collaboration is
not ready to share their RFQD, unless this would be the last resort
for keeping the AD program alive.
R.Landua is investigating the possibility of funding
via the European Framework program 6 (improvement of existing
research facilities). Is there any sense in building a second RFQD since
the space in the AD hall is limited?
- PS area modifications during the next shut down. Following the
proposal M. Chanel presented at the last PSMB for modifications in the PS
area during the next shut down, a planning for the installation work has been
prepared. The work will start as soon as possible in January 2003 and be finished
on 10 March 2003, depending on different decisions on the availability of
magnets and mechanical parts.
- ABMB. J.P. Delahaye gave a summary of the meeting held on 21.10.2002
LHC financial and technical status report has been completed.
- Staff Plan : the staff plan data from AB, AC and AT were combined
and presented by C. Wyss to the DG on 17 October 2002.
- A&B budget codes : A proposal for the new codes was presented
- The structure of groups/sections in the new A&B Division
will be presented in a meeting which will take place on Thursday 31.10.200
at 2 pm.
- Common Control Room : A meeting took place last week with R.
Bailey, S. Baird, C. Wyss and the architect to discuss the building and
design. For future meetings a representative of the TCR should also be
- "Chamonix 2003" : will take place at Chamonix from
3 to 8 March 2003.
- Space Needs: A global plan for the whole A&B division will
be discussed in a future meeting.
- CERN Management Board : K. Schindl attended the CERN Management Board
which took place on Thursday 24.10.2002. The following points were discussed
(see summary of
K. Schindl) :
- Tref Meeting
- Preparation of the Finance Committee 6th November 2003
- LHC Review Panel
- Preparation Morges III
- Appointment of the Deputy Director General : R. Cashmore
- the TEACC and CP changes results were presented and discussed in a Closed
- Beryllium inventory : Following a request from TIS (IS 25), G. Daems
will have to list all beryllium used in the PS division. The inventory must
show chemical composition (e.g. oxide or metal), form, weight and where used
November 2002 : Group Leaders should send their requests and prolongations
proposals by Wednesday 30 October 2002, 12:00 to Tjitske Kehrer.
- Shut down : After the last machine development which took place on Monday
morning 28.10.2002, the PS Complex was shut down at 14:00.
- Operation Statistics : S. Baird will present the 2002 statistics
of operation in a future PSMB Meeting this year.
| PS Division | PSMB
Meetings | Other PS Meetings |
29 October, 2002